OpenStudy (anonymous):

Use the substitution method to find the solution(s) to the system of equations. Show all work for full credit. y = x2 y = x + 2

OpenStudy (anonymous):

\[ \begin{align} y&=x^2\\ y&=x+2 \end{align} \] Where are you stuck..?

OpenStudy (anonymous):

I'm not sure how to solve it. I keep getting no solution.

OpenStudy (matheducatormcg):

you need to make a substitution so that you are relating the x variables to eachother.

OpenStudy (anonymous):

Put \(x^2\) where \(y\) is in \(y=x+2.\)

OpenStudy (anonymous):

Solve that equation. Then find \(y\).

OpenStudy (anonymous):

Y=X2 X2=X+2 X2-X-2=0 (X+1)(X-2)=0 Continue from there Find X & replace on the equation to find Y Simple

OpenStudy (matheducatormcg):

you could also graph both equations and look for the point of intersection.

OpenStudy (anonymous):

ThanksI

OpenStudy (anonymous):

You're welcome! Have a good day now.